# AutoEP: LLMs-Driven Automation of Hyperparameter Evolution for Metaheuristic Algorithms

### v1

@@ Abstract
Dynamically configuring algorithm hyperparameters is a fundamental challenge in computational intelligence. While learning-based methods offer automation, they suffer from prohibitive sample complexity and poor generalization. We introduce AutoEP, a novel framework that bypasses training entirely by leveraging Large Language Models (LLMs) as zero-shot reasoning engines for algorithm control. AutoEP’s core innovation lies in a tight synergy between two components: (1) an online Exploratory Landscape Analysis (ELA) module that provides real-time, quantitative feedback on the search dynamics, and (2) a multi-LLM reasoning chain that interprets this feedback to generate adaptive hyperparameter strategies. This approach grounds high-level reasoning in empirical data, mitigating hallucination. Evaluated on three distinct metaheuristics across diverse combinatorial optimization benchmarks, AutoEP consistently outperforms state-of-the-art tuners, including neural evolution and other LLM-based methods. Notably, our framework enables open-source models like Qwen3-30B to match the performance of GPT-4, demonstrating a powerful and accessible new paradigm for automated hyperparameter design. Our code is available at https://anonymous.4open.science/r/AutoEP-3E11.
@@ 1 Introduction
The performance of complex algorithms, from numerical optimizers to machine learning models, is critically governed by their internal hyperparameters. Dynamically adapting these parameters to suit the problem instance at hand represents a long-standing challenge in automated algorithm design Wu et al. (2024; 2025). Metaheuristic algorithms, a cornerstone of solving complex combinatorial problems, serve as a canonical example. Their effectiveness hinges on a delicate balance between exploration (diversifying the search) and exploitation (intensifying the search in promising regions), a trade-off directly controlled by their hyperparameter configurations Eiben et al. (2002a). Mastering this balance is crucial for achieving state-of-the-art performance but remains a formidable open problem.
Traditional approaches to dynamic hyperparameter adaptation fall into two categories. Manual, rule-based strategies Eiben et al. (2002b); Thierens (2005); Ansótegui et al. (2009); Joshi & Bansal (2020a); Leon & Xiong (2015); Li et al. (2025); Shao et al. (2025) embed human expertise into hard-coded logic, but are brittle, labor-intensive, and fail to generalize across different problems or algorithms. To overcome this, data-driven methods, particularly deep reinforcement learning (DRL) Tessari & Iacca (2022); Ma et al. (2024); Liu et al. (2023); Tatsis & Parsopoulos (2020); Yin et al. (2021), have attempted to learn adaptive policies from scratch. However, this paradigm faces fundamental limitations: (1) prohibitive sample complexity, requiring millions of algorithm executions to train a single policy, and (2) poor generalization, where policies often overfit to the training distribution of problems and fail on unseen instances or algorithm variants Guo et al. (2024). This reveals a critical gap: the need for a framework that can adapt algorithm behavior without requiring expensive, instance-specific training.
@@
(4) State-of-the-Art Performance Across Diverse Benchmarks: Through extensive experiments on three distinct metaheuristics (GA, PSO, ACO Holland (1992); Dorigo et al. (2007); Kennedy & Eberhart (1995)) and four combinatorial optimization problems, we show that AutoEP consistently and significantly outperforms both traditional hyperparameter tuning methods and recent LLM-based approaches.
@@ 2 Related work
Rule-Based and Heuristic Control. Early attempts to automate hyperparameter tuning relied on hard-coded, rule-based heuristics. These methods embed expert knowledge into predefined rules that adjust parameters based on simple metrics like iteration count or population diversity Thierens (2005). For instance, strategies might deterministically increase mutation rates to escape local optima or adapt selection pressure over time Joshi & Bansal (2020a). While an improvement over static settings, these approaches are fundamentally brittle. The underlying heuristics are problem-dependent and require extensive manual calibration. They lack the ability to adapt to unforeseen dynamics in the search process, making them unable to generalize across different problem classes or algorithms.
Learning Control Policies from Scratch via Reinforcement Learning. To overcome the rigidity of manual heuristics, a significant body of work has focused on learning control policies from scratch using Reinforcement Learning (RL) Ma et al. (2024); Liu et al. (2023); Tatsis & Parsopoulos (2020). In this paradigm, a metaheuristic’s state (e.g., population statistics) is used by an RL agent to select an action (i.e., a new hyperparameter configuration), receiving a reward based on the improvement in solution quality. Frameworks like GLEET Ma et al. (2024) and NeuroCrossover Liu et al. (2023) have successfully used policy gradient algorithms like PPO to train agents for specific evolutionary operators or parameters. Despite their success, these methods suffer from the characteristic limitations of RL: prohibitive sample complexity. They require millions of computationally expensive simulation runs to converge to a robust policy, and the resulting agent often overfits to the specific algorithm and problem distribution it was trained on, exhibiting poor generalization to new scenarios Guo et al. (2024).
@@
LLMs for Algorithm Design: From Offline Generation to Online Control. The advent of LLMs has introduced a new frontier for algorithm design, leveraging their vast pre-trained knowledge of code and logic. Seminal works like EoH Liu et al. (2024) and ReEvo YE et al. (2024) have demonstrated that LLMs can act as powerful offline algorithm generators. They can be prompted to create novel heuristic operators or suggest entire initial configurations, effectively automating parts of the algorithm design process. However, these approaches are predominantly static: the LLM is consulted before the optimization run to generate code or a configuration, but it does not participate in the process itself. This leaves a critical gap, as the optimal strategy for a metaheuristic is not static but state-dependent, evolving as the search progresses Surina et al. (2025). The true opportunity lies in using an LLM not as an offline code generator, but as a real-time online reasoning engine that dynamically steers the algorithm based on live feedback. Our work is the first to address this gap, proposing a framework for online, dynamic hyperparameter control.
In summary, prior work has established the need for dynamic hyperparameter control but has fallen short of a generalizable, training-free solution. Rule-based methods are brittle, RL-based methods are sample-inefficient, and current LLM applications remain confined to offline generation. This highlights the clear need for a framework that enables online, adaptive control by grounding the reasoning capabilities of LLMs in the real-time dynamics of the search process.
@@ 3.1 Grounding Reasoning with Quantitative Search Dynamics
To effectively control a metaheuristic algorithm, the decision-making agent requires a real-time, quantitative understanding of the search process. As metaheuristics are black-box methods, we employ Exploratory Landscape Analysis (ELA) Mersmann et al. (2011); Ma et al. (2025) to extract features that characterize the algorithm’s state. We selected a concise yet comprehensive set of features designed to capture four key aspects of the search: (1) the statistical distribution of the current population’s fitness, (2) the structural properties of the local fitness landscape, (3) the diversity of solutions, and (4) the recent progress of the search.
@@ 4.1 Experimental settings
To comprehensively evaluate AutoEP’s performance, we selected classic combinatorial optimization datasets, including TSP Matai et al. (2010), CVRP Dantzig & Ramser (1959), and FSSPEmmons & Vairaktarakis (2012) , as well as a more complex, realistic optimization task involving UAV-enabled IoT data collection Zhan et al. (2017). We compared AutoEP against three categories of algorithms:
Hyperparameter tuning methods for metaheuristic algorithms. PT Joshi & Bansal (2020b) is a recent manually designed method for hyperparameter tuning. GLEET Ma et al. (2024) represents the SOTA in reinforcement learning-based hyperparameter tuning; we retrained the network strictly following the original experimental settings for different algorithms and datasets. BEA Lan et al. (2022) is a leading method using bayesian optimization for hyperparameter tuning.
@@ 4.2.1 Validation on TSP
For the TSP problem, we selected the TSPLIBReinelt (1991) dataset and used GAHolland (1992) and GA-2opt Sabba & Chikhi (2013) as baseline algorithms. GA is a widely used metaheuristic algorithm, while GA-2opt, which combines global and local search, is a robust heuristic for TSP. Detailed experimental results are presented in Table 1.The first row displays the performance of neural combinatorial optimization methods on various TSP datasets. The second row compares hyperparameter tuning methods and LLMs-enhanced metaheuristic operators using the GA algorithm. Among hyperparameter tuning methods, GLEET performed the best, while ReEvo showed promising results in enhancing metaheuristic operators. AutoEP, after optimizing GA’s hyperparameters, achieved the best results across all problem sizes.The third row evaluates GA-2opt, which combines population-based and local search strategies, resulting in strong performance. The test results indicate that the algorithm with dynamically controlled hyperparameters by AutoEP achieved SOTA results across all test datasets, surpassing current neural combinatorial optimization SOTA methods like LEHD and DACT. These comparative results demonstrate that AutoEP, as a plug-and-play framework for tuning metaheuristic algorithm hyperparameters, significantly enhances the performance of the original algorithms. To validate AutoEP’s ability to dynamically adjust hyperparameters when integrated as a plugin with any metaheuristic algorithm, we applied it to ReEvo and EoH-enhanced GA-2opt algorithms, as shown in the fourth row. Our results demonstrate that AutoEP further improves the performance of these enhanced algorithms, with final results closely matching those of GA-2opt+AutoEP. This confirms two key points:
@@ 5 Discussion and Conclusion
Figure 5: Comparison of hyperparameter tuning frequencies.
Practical Advantages for Real-World Optimization. Beyond its novelty, AutoEP is designed for practicality. Its plug-and-play framework makes it a general-purpose tool for enhancing any metaheuristic algorithm. Crucially, as demonstrated by our ablation studies, AutoEP’s structured reasoning framework reduces the dependency on a single, monolithic LLM’s raw intelligence. This architectural strength allows it to achieve SOTA performance using smaller, open-source models (e.g., 30B-72B class). This is a critical advantage for real-world applications like factory scheduling or logistics, where local deployment is necessary to ensure data privacy, low latency, and operational reliability, and where deploying massive proprietary models is often infeasible.
Appendix A Technical Reproducibility and LLM Usage
Our paper has no conflicts of interest and complies with ethical standards. Our paper code is reproducible, and we have provided an anonymous link to the reproducible code. We only used LLMs to polish the paper writing.
@@ B.1 Validation on CVRP
For the CVRP problem, we used the VRPLIB Semet & Taillard (1993) dataset and included both GA-2opt and PSO-2opt Sabba & Chikhi (2013) algorithms to evaluate AutoEP’s performance across different metaheuristic algorithms. The test results are presented in Table 4.In the first column, DACT and LEHD continue to show strong performance. The second column compares various methods for improving PSO-2opt, where AutoEP achieves superior results. The third column evaluates enhancements to GA-2opt, with AutoEP demonstrating the best performance. Additionally, GA-2opt combined with AutoEP achieves the smallest gap from the optimal solutions across all datasets.
@@ B.2 Validation on FSSP
For the FSSP problem, we used the Taillard Taillard (1993) dataset and GA-2opt as the baseline algorithm. We also included advanced methods for FSSP: NEH Nawaz et al. (1983), NEHFF Fernandez-Viagas & Framinan (2014), and PFSPNet_NEH Pan et al. (2021). The test results are presented in Table 5. In the first column, PFSPNet_NEH shows superior performance among the comparison algorithms. The second column demonstrates that AutoEP significantly enhances GA-2opt across all datasets, consistently achieving the best results.

### v2

@@ Abstract
Dynamically configuring algorithm hyperparameters is a fundamental challenge in computational intelligence. While learning-based methods offer automation, they suffer from prohibitive sample complexity and poor generalization. We introduce AutoEP, a novel framework that bypasses training entirely by leveraging Large Language Models (LLMs) as zero-shot reasoning engines for algorithm control. AutoEP’s core innovation lies in a tight synergy between two components: (1) an online Exploratory Landscape Analysis (ELA) module that provides real-time, quantitative feedback on the search dynamics, and (2) a multi-LLM reasoning chain that interprets this feedback to generate adaptive hyperparameter strategies. This approach grounds high-level reasoning in empirical data, mitigating hallucination. Evaluated on three distinct metaheuristics across diverse combinatorial optimization benchmarks, AutoEP consistently outperforms state-of-the-art tuners, including neural evolution and other LLM-based methods. Notably, our framework enables open-source models like Qwen3-30B to match the performance of GPT-4, demonstrating a powerful and accessible new paradigm for automated hyperparameter design. Our code is available at https://github.com/YiZheZhang12/AutoEP.
@@ 1 Introduction
The performance of complex algorithms, from numerical optimizers to machine learning models, is critically governed by their internal hyperparameters. Dynamically adapting these parameters to suit the problem instance at hand represents a long-standing challenge in automated algorithm design Wu et al. (2024; 2025). Metaheuristic algorithms, a cornerstone of solving complex combinatorial problems, serve as a canonical example. Their effectiveness hinges on a delicate balance between exploration (diversifying the search) and exploitation (intensifying the search in promising regions), a trade-off directly controlled by their hyperparameter configurations Eiben et al. (2002). Mastering this balance is crucial for achieving state-of-the-art performance but remains a formidable open problem.
Traditional approaches to dynamic hyperparameter adaptation fall into two categories. Manual, rule-based strategies Eiben et al. (2002); Thierens (2005); Ansótegui et al. (2009); Joshi and Bansal (2020); Leon and Xiong (2015); Li et al. (2025b); Shao et al. (2025) embed human expertise into hard-coded logic, but are brittle, labor-intensive, and fail to generalize across different problems or algorithms. To overcome this, data-driven methods, particularly deep reinforcement learning (DRL) Tessari and Iacca (2022); Ma et al. (2024); Liu et al. (2023); Tatsis and Parsopoulos (2020); Yin et al. (2021), have attempted to learn adaptive policies from scratch. However, this paradigm faces fundamental limitations: (1) prohibitive sample complexity, requiring millions of algorithm executions to train a single policy, and (2) poor generalization, where policies often overfit to the training distribution of problems and fail on unseen instances or algorithm variants Guo et al. (2024). This reveals a critical gap: the need for a framework that can adapt algorithm behavior without requiring expensive, instance-specific training.
@@
(4) State-of-the-Art Performance Across Diverse Benchmarks: Through extensive experiments on three distinct metaheuristics (GA, PSO, ACO Holland (1992); Dorigo et al. (2007); Kennedy and Eberhart (1995)) and four combinatorial optimization problems, we show that AutoEP consistently and significantly outperforms both traditional hyperparameter tuning methods and recent LLM-based approaches.
@@ 2 Related work
Rule-Based and Heuristic Control. Early attempts to automate hyperparameter tuning relied on hard-coded, rule-based heuristics. These methods embed expert knowledge into predefined rules that adjust parameters based on simple metrics like iteration count or population diversity Thierens (2005). For instance, strategies might deterministically increase mutation rates to escape local optima or adapt selection pressure over time Joshi and Bansal (2020). While an improvement over static settings, these approaches are fundamentally brittle. The underlying heuristics are problem-dependent and require extensive manual calibration. They lack the ability to adapt to unforeseen dynamics in the search process, making them unable to generalize across different problem classes or algorithms.
Meta-Black-Box Optimization (Meta-BBO). To address the inherent rigidity of manual heuristics, the field has shifted toward Meta-BBO Ma et al. (2023); Tatsis and Parsopoulos (2020); Guo et al. (2025a); Ma et al. (2025c); Li et al. (2025a). This paradigm integrates data-driven and reinforcement learning (RL) approaches to automate algorithm configuration and design. Unlike traditional optimization that searches within a decision space, Meta-BBO explores the algorithm space to discover optimal optimizers, typically leveraging historical data to learn adaptive strategies. While early frameworks like GLEET Ma et al. (2024) and NeuroCrossover Liu et al. (2023) successfully employed Deep RL for operator selection, they were constrained by high sample complexity. Recent advancements have expanded this domain; for instance, Neural ELA Ma et al. (2025a) incorporates landscape features via neural exploratory landscape analysis, while DesignX Guo et al. (2025b) proposes a dual-agent RL framework to jointly learn algorithm structure and hyperparameters. Despite these advances, current Meta-BBO methodologies remain heavily reliant on computationally expensive meta-training. This highlights a critical gap: the need for a zero-shot framework that retains the adaptivity of Meta-BBO while circumventing prohibitive training costs.
LLMs for Algorithm Design: From Offline Generation to Online Control. Leveraging pre-trained code knowledge, LLMs have advanced algorithm design significantly Ma et al. (2025b). Offline frameworks (e.g., EoH Liu et al. (2024), ReEvo YE et al. (2024)) successfully automate the generation of operators and configurations. However, optimization is inherently dynamic Surina et al. (2025), necessitating online adaptation. Recent attempts like EvoLLM Lange et al. (2024) employ LLMs as evolutionary operators for direct solution generation. Despite their novelty, such methods struggle with floating-point representation and context window constraints in high-dimensional search spaces.
This highlights a critical research gap. We argue that the optimal role for an LLM is not to mimic the solver, but to act as a high-level supervisor. Unlike prior works that replace search operators, we propose a framework for dynamic hyperparameter control. This design grounds the LLM’s semantic reasoning in real-time search dynamics without being hindered by numerical precision issues. By doing so, we overcome the brittleness of rule-based heuristics and the sample inefficiency of RL, offering a robust and scalable solution for online control.
@@ 3.1 Grounding Reasoning with Quantitative Search Dynamics
To effectively control a metaheuristic algorithm, the decision-making agent requires a real-time, quantitative understanding of the search process. As metaheuristics are black-box methods, we employ Exploratory Landscape Analysis (ELA) Mersmann et al. (2011); Ma et al. (2025a) to extract features that characterize the algorithm’s state. We selected a concise yet comprehensive set of features designed to capture four key aspects of the search: (1) the statistical distribution of the current population’s fitness, (2) the structural properties of the local fitness landscape, (3) the diversity of solutions, and (4) the recent progress of the search.
@@ 4.1 Experimental settings
To comprehensively evaluate AutoEP’s performance, we selected classic combinatorial optimization datasets, including TSP Matai et al. (2010), CVRP Dantzig and Ramser (1959), and FSSP Emmons and Vairaktarakis (2012), as well as a more complex, realistic optimization task involving UAV-enabled IoT data collection Zhan et al. (2017). We compared AutoEP against three categories of algorithms:
Hyperparameter tuning methods for metaheuristic algorithms. PT Joshi and Bansal (2020) is a recent manually designed method for hyperparameter tuning. GLEET Ma et al. (2024) represents the SOTA in reinforcement learning-based hyperparameter tuning; we retrained the network strictly following the original experimental settings for different algorithms and datasets. BEA Lan et al. (2022) is a leading method using bayesian optimization for hyperparameter tuning.
@@ 4.2.1 Validation on TSP
For the TSP problem, we selected the TSPLIBReinelt (1991) dataset and used GAHolland (1992) and GA-2opt Sabba and Chikhi (2013) as baseline algorithms. GA is a widely used metaheuristic algorithm, while GA-2opt, which combines global and local search, is a robust heuristic for TSP. Detailed experimental results are presented in Table 1.The first row displays the performance of neural combinatorial optimization methods on various TSP datasets. The second row compares hyperparameter tuning methods and LLMs-enhanced metaheuristic operators using the GA algorithm. Among hyperparameter tuning methods, GLEET performed the best, while ReEvo showed promising results in enhancing metaheuristic operators. AutoEP, after optimizing GA’s hyperparameters, achieved the best results across all problem sizes.The third row evaluates GA-2opt, which combines population-based and local search strategies, resulting in strong performance. The test results indicate that the algorithm with dynamically controlled hyperparameters by AutoEP achieved SOTA results across all test datasets, surpassing current neural combinatorial optimization SOTA methods like LEHD and DACT. These comparative results demonstrate that AutoEP, as a plug-and-play framework for tuning metaheuristic algorithm hyperparameters, significantly enhances the performance of the original algorithms. To validate AutoEP’s ability to dynamically adjust hyperparameters when integrated as a plugin with any metaheuristic algorithm, we applied it to ReEvo and EoH-enhanced GA-2opt algorithms, as shown in the fourth row. Our results demonstrate that AutoEP further improves the performance of these enhanced algorithms, with final results closely matching those of GA-2opt+AutoEP. This confirms two key points:
@@ 4.5 Sensitivity to Adjustment Frequency
A critical concern in using LLMs for online control is the management of context length. As noted in Sec3.2, the Experience Pool stores historical states and actions. To prevent prompt bloat and minimize context noise, AutoEP employs a sliding window mechanism, retaining only the most recent $L$ iterations. We conducted an ablation study on the TSP-100 dataset using GA-2opt to determine the optimal window size. We compared short ($L=5$), medium ($L=20$), long ($L=50$), and infinite (Full History) window sizes.
Figure 5: Comparison of hyperparameter tuning frequencies(UAV-300).
As shown in Table 4, utilizing the full history significantly degrades both efficiency and solution quality. The performance drop in the "Full History" setting confirms that irrelevant historical data (e.g., early-stage exploration metrics) acts as context noise during later optimization stages, causing the LLM to generate suboptimal strategies. Conversely, a very small window ($L=5$) is computationally efficient but lacks sufficient temporal context to identify complex trends like stagnation. The default setting of $L=20$ strikes the optimal balance, providing sufficient history for trend analysis while maintaining low inference latency and filtering out obsolete data.
Table 4: Impact of experience pool size ($L$) on performance and inference latency (TSP-100).
Pool Size ($L$) Opt. Gap (%) Avg. Inference Latency (s) Observation
$L=5$ 0.04 0.18 Limited context; reactive behavior
$L=20$ (Default) 0.01 0.31 Balanced context and efficiency
$L=50$ 0.03 0.85 Increased latency; minor distractions
Full History 0.17 >2.50>2.50 High latency; hallucination from noise
4.6 Interpretability Analysis: Visualizing LLM-Driven Control Dynamics
To demonstrate the interpretability of AutoEP, we visualize the evolutionary trajectory of GA hyperparameters (Mutation and Crossover probabilities) on the TSP-400 instance (Figure 6). The resulting curves reveal a sophisticated, state-aware control strategy that distinctively contrasts with manual, monotonic schedules.
Response to ELA Features (Attention to State): The LLM’s decisions are explainable through the observed ELA metrics. Periods of rapid parameter fluctuation correspond to specific landscape characteristics. For example, when the Dispersion Ratio ($D_{ratio}$) indicates a single-funnel structure4, AutoEP prioritizes exploitation. Conversely, when Skewness signals a risk of premature convergence5, the system triggers a ’rescue’ behavior by sharply increasing the mutation rate.
Figure 6: Visualization of hyperparameter evolution for GA on TSP-400.
Operationalizing the Exploration-Exploitation Trade-off: The visualization confirms that AutoEP understands the mechanics of GA. The crossover and mutation curves often mirror each other (inverse correlation). This validates that the CoR engine successfully translates abstract directives (e.g., ’intensify exploration’) into logically consistent numerical actions (High Mutation / Low Crossover).
Emergence of Search Policies via In-Context Learning: Beyond direction, the magnitude of adjustments shows structured patterns. The LLM utilizes the Experience Pool to determine not just what to change, but how much. This results in distinct phases: aggressive large-step adjustments to escape stagnation, followed by fine-grained small-step tuning for local convergence, proving the effective utilization of historical search context.
@@ 5 Discussion and Conclusion
Practical Advantages for Real-World Optimization. Beyond its novelty, AutoEP is designed for practicality. Its plug-and-play framework makes it a general-purpose tool for enhancing any metaheuristic algorithm. Crucially, as demonstrated by our ablation studies, AutoEP’s structured reasoning framework reduces the dependency on a single, monolithic LLM’s raw intelligence. This architectural strength allows it to achieve SOTA performance using smaller, open-source models (e.g., 30B-32B class). This is a critical advantage for real-world applications like factory scheduling or logistics, where local deployment is necessary to ensure data privacy, low latency, and operational reliability, and where deploying massive proprietary models is often infeasible.
Ethical Statement and Reproducibility Statement
Our paper has no conflicts of interest and complies with ethical standards. Our paper code is reproducible, and we have provided an anonymous link to the reproducible code.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant Nos. 62273352, 72501290, and 72401287. The authors would like to thank the anonymous reviewers for their constructive feedback, which helped improve the quality of this paper.
Appendix A LLM Usage
We only used LLMs to polish the paper writing.
@@ B.1 Validation on CVRP
For the CVRP problem, we used the VRPLIB Semet and Taillard (1993) dataset and included both GA-2opt and PSO-2opt Sabba and Chikhi (2013) algorithms to evaluate AutoEP’s performance across different metaheuristic algorithms. The test results are presented in Table 5.In the first column, DACT and LEHD continue to show strong performance. The second column compares various methods for improving PSO-2opt, where AutoEP achieves superior results. The third column evaluates enhancements to GA-2opt, with AutoEP demonstrating the best performance. Additionally, GA-2opt combined with AutoEP achieves the smallest gap from the optimal solutions across all datasets.
@@ B.2 Validation on FSSP
For the FSSP problem, we used the Taillard Taillard (1993) dataset and GA-2opt as the baseline algorithm. We also included advanced methods for FSSP: NEH Nawaz et al. (1983), NEHFF Fernandez-Viagas and Framinan (2014), and PFSPNet_NEH Pan et al. (2021). The test results are presented in Table 6. In the first column, PFSPNet_NEH shows superior performance among the comparison algorithms. The second column demonstrates that AutoEP significantly enhances GA-2opt across all datasets, consistently achieving the best results.